Beth is twice as old as her brother and half as old as her father. in 22 years, her brother will be half as old as his father, how old is beth now
Answer by richwmiller(17219)

b=Beth now
r=brother now
f=father now
b=2r
b/2=r
b=f/2
(r+22)=(f+22)/2
b/2+22=b+11
b/2+11=b
11=b/2
22=b
r=11
f=44
Beth is 22 now.